# An AI Implementation Science Study to Improve Trustworthy Data in a Large Healthcare System

### v1

@@ Abstract
Preprint Notice: This is the author’s preprint version of a manuscript accepted at IEEE BHI 2025. The final version will appear in IEEE Xplore.
I Introduction
@@
II Background
II-A SC Data Infrastructure
@@ II-D METRIC Framework
II-E Case Study: Craniofacial Microsomia
@@ III-A Modernization and Evaluation of SC RDW
III-A1 AI Implementation Challenges and Opportunities
@@
III-B TAI-based Implementation Evaluation
@@
III-C Systematic vs. Case Study
@@
IV Results
IV-A Improving SC Data Infrastructure
@@
IV-B Implementation Evaluation Analysis
@@
TABLE I: Overlapping between ICD9-ICD10 and OMOP CDM codes
For RQ3 (Timeliness), we compared the overlap between ICD-9 and ICD-10 procedure codes. We found that only half of the ICD-9 codes shared a common mapping with an ICD-10 code (Table I). This limited overlap suggests that AI models are at risk of performance degradation when encountering data with different distributions of these coding systems. While ICD-10 codes were more prevalent, likely due to their greater comprehensiveness [topaz2013icd], the poor mapping indicates a broader potential risk that may affect other clinical vocabularies and warrants further investigation.
@@
IV-C Case Study Specificity
@@ IV-C1 Data Retrieval
IV-C2 Data Fidelity and Usefulness
@@
IV-C3 CFM Case-Study and FHIR Implementation
@@ VI Conclusion
The CFM case study involving human subjects was conducted in accordance with the ethical standards outlined in the Belmont Report and received approval from the [REDACTED], IRB approval number [REDACTED].

### v2

@@ Abstract
Keywords: health informatics, FHIR, OMOP-CDM, data standard, data harmonization, data quality, trustworthy AI, AI implementation science
1 Introduction
@@
2 Background
2.1 SC Data Infrastructure
@@ 2.4 METRIC Framework
2.5 Case Study: Craniofacial Microsomia
@@ 3.1 Modernization and Evaluation of SC RDW
3.1.1 AI Implementation Challenges and Opportunities
@@
3.2 TAI-based Implementation Evaluation
@@
3.3 Systematic vs. Case Study
@@
4 Results
4.1 Improving SC Data Infrastructure
@@
4.2 Implementation Evaluation Analysis
@@
Table 1: Overlapping between ICD9-ICD10 and OMOP CDM codes
For RQ3 (Timeliness), we compared the overlap between ICD-9 and ICD-10 procedure codes. We found that only half of the ICD-9 codes shared a common mapping with an ICD-10 code (Table 1). This limited overlap suggests that AI models are at risk of performance degradation when encountering data with different distributions of these coding systems. While ICD-10 codes were more prevalent, likely due to their greater comprehensiveness [topaz2013icd], the poor mapping indicates a broader potential risk that may affect other clinical vocabularies and warrants further investigation.
@@
4.3 Case Study Specificity
@@ 4.3.1 Data Retrieval
4.3.2 Data Fidelity and Usefulness
@@
4.3.3 CFM Case-Study and FHIR Implementation
@@ 6 Conclusion
The CFM case study involving human subjects was conducted in accordance with the ethical standards outlined in the Belmont Report and received approval from the Georgia Institute of Technology, IRB approval number H21297.